4. Integration: TT (a) Si Jachtvoz dx (b) sin x dx (6-cos x)3 (c) , (2-2) (3) dx (d) St(t - 5)8dt 6x7-x*+VX-4 dx x2
Answer #1

solution 4(a) lit I= 9 앉 1HJ2 다 at itsnet 어스가 d2d ㄴ ㅗ 아 = J전 하 29 터 145 5 -- 113 t = 4 와며 스 - IN 2 4. 2dt 43 2. lusing = tra= الم N IN { Ans 었 () 싸 SR 앉 3. (6 - let 아 6 4 sinredna 아 It n=0 스 6 Go =6-1 하 5 = 6 건 fin Sind 2 3 درنا - 6 5 5 11 -3 t dt2. ( 1 3 2. (0) * ( (= - == ما مرا 7 A ( - ) ر - ال : ہر 3 - 2 dna dr فير ۔ ر = I ہو ( - )) * 2 2 اے / w * الى - 2 )7 ر2 ہا ل② let I-st lt-500 dt * = ك - dt de م ل-2 دو هد ( 5 / - 1 - flor ره ( 2 5 7 ۶ / ہ - 5*۶ = / و % 10 9 - (۴۶) کی مرد) (t-s) يح A